# VoiceBridge: Designing Latent Bridge Models for General Speech Restoration at Scale

### v1

@@ F.1 Evaluation datasets
As GSR is defined as general restoration, including different restoration processes, the restoration process from a single degradation naturally becomes its subtask. To evaluate VoiceBridge’s zero-shot performance on these restoration subtasks, we choose the Speech Enhancement (SE) task as a typical example. We utilize VoiceBank-Demand (Valentini-Botinhao, 2017) and WSJ0-CHiME3 (Barker et al., 2015). The VoiceBank-Demand dataset mixes utterances from the Voicebank speech corpus with real-world noise from the DEMAND dataset, serving as a benchmark testset for the SE. WSJ0-CHiME3, similarly, mixes clean speech signals from the WSJ0 dataset with noise from CHiME3, but with a smaller signal-to-noise ratio, making the denoising task harder. The two benchmark datasets are widely used in prior SE works (Lu et al., 2021; 2022; Richter et al., 2023; Lemercier et al., 2023b; Jukić et al., 2024).
@@
For the CAR test, we utilize the VCTK (Veaux & Yamagishi, 2017) testset, using Encodec (Défossez et al., 2022) to compress the 48kHz audio to a codec at a 3kbps frame rate, and then reconstruct it.
@@
Table 10: Denoising results on VCTK (Veaux & Yamagishi, 2017) with additive noise from DEMAND (Thiemann et al., 2013) at different SNR levels.
Dataset Method PESQ($\uparrow$) WVMOS($\uparrow$) UTMOS($\uparrow$) DNSMOS($\uparrow$) NISQA($\uparrow$) SNR=-15dB SBSE 1.119 -0.949 1.480 1.513 1.869 Voicefixer 1.225 2.423 2.278 2.313 2.848 Resemble-Enhance 1.168 3.394 2.440 2.640 3.130 UniverSE++ 1.211 2.772 2.271 2.374 2.950 VoiceBrigde 1.352 3.479 2.755 2.698 3.270 SNR=-10dB SBSE 1.224 0.347 1.829 1.822 2.389 Voicefixer 1.588 3.263 2.845 2.791 3.576 Resemble-Enhance 1.345 3.783 2.929 2.876 3.693 UniverSE++ 1.527 3.704 3.075 2.933 3.971 VoiceBrigde 1.453 3.908 3.339 2.944 3.725 SNR=-5dB SBSE 1.410 1.698 2.301 2.275 2.933 Voicefixer 1.871 3.901 3.346 3.106 4.133 Resemble-Enhance 1.586 4.001 3.266 3.009 4.075 UniverSE++ 1.790 3.820 3.320 2.913 4.177 VoiceBrigde 1.730 4.093 3.641 3.022 3.969 SNR=0dB SBSE 1.663 2.938 2.826 2.665 3.468 Voicefixer 2.143 4.074 3.548 3.170 4.350 Resemble-Enhance 1.891 4.135 3.491 3.070 4.322 UniverSE++ 2.240 4.100 3.674 2.974 4.291 VoiceBrigde 2.200 4.200 3.765 3.098 4.201
@@
VoiceBridge 3.45 3.99 3.17 4.50
F.2 Baseline Methods
F.3 Metrics
F.4 Degradation Breakdown
To make comprehensive evaluation, we include task-specific baseline models and evaluation metrics. For the BWE task, we include audio super-resolution models and other speech BWE model, including AudioSR (Liu et al., 2024), NuWave2 (Han & Lee, 2022), TFiLM (Birnbaum et al., 2019), and HiFi++ (Andreev et al., 2023). For dereverberation on WSJ-Reverb, we include SGMSE+ (Richter et al., 2023; Welker et al., 2022) and StoRM (Lemercier et al., 2023b). For SE, we include SBSE (Jukić et al., 2024). Additional metrics involves the following: STOI (Taal et al., 2010) is an objective metric that measures the intelligibility of degraded speech. LSD (Liu et al., 2022) directly calculates the log-spectrogram-distance between the enhanced and ground truth signals, which is a commonly used metric in BWE scenarios (Liu et al., 2024; Li et al., 2025a). SRMR (Falk et al., 2010), a non-intrusive metric specially designed for measuring signal to reverberation energy ratio, is utilized for measuring dereverberation performance in our experiments.
F.5 Comparison with Pretrained Models
@@
Appendix G Further Studies on Latent Modeling
G.1 Quantitative Results on the Joint Neural Prior
To quantitatively analyze the effect of prior convergence beyond tSNE visualization, we calculate the 2-Wasserstein distance (Panaretos & Zemel, 2019) matrix over the latent distribution for different priors. We model the latent distribution for each degradation type as a Gaussian distribution with the sample mean and variance. The 2-Wasserstein (Fréchet) distance between two Gaussian-fitted distributions $\mathcal{N}(\mu_{i},\Sigma_{i})$ and $\mathcal{N}(\mu_{j},\Sigma_{j})$ is defined as
@@ G.1 Quantitative Results on the Joint Neural Prior
G.2 Exceeding the Perceptual Quality Limits of Latent Models
Without our proposed perceptual-aware fine-tuning process, the LBM is trained to capture the distribution of the latent of the HQ target, which may not be fully aligned with human perceptual quality. Moreover, without joint fine-tuning, the perceptual generation quality of VoiceBridge will be limited by VAE as well. As the perceptual-aware objectives have not been explicitly considered in the training of VAE, the perceptual quality of VoiceBridge will be inevitably restricted. By the perceptual-aware joint fine-tuning process, we can improve the perceptual generation quality of both LBMs and the VAE decoder, meanwhile reducing their cascading error in generation.
@@ G.2 Exceeding the Perceptual Quality Limits of Latent Models
To verify that VoiceBridge surpasses the upper-bound reconstruction performance of the original VAE, we compare the restoration results of noisy inputs with EP-VAE reconstructions of clean ones. Specifically, we report four groups of results: (1) the EP-VAE reconstruction of clean audio (VAE-Rec.); (2) the restoration results of corresponding degraded audio of VoiceBridge (VB-Gen.); (3) the restoration results of VoiceBridge when passed the clean audio as input (VB-Rec.), which can be seen as a version of reconstruction with enhanced perceptual quality by VoiceBridge; (4) the ground-truth clean target (GT). Clean and degraded data are taken from the VoiceBank-Demand (Valentini-Botinhao, 2017) dataset. As shown in Table 12, the perceptual-aware fine-training stage enables VoiceBridge’s restoration to surpass the VAE reconstruction limits in perceptual quality. Moreover, when GT is passed as input for VoiceBridge instead of its noisy version, the perceptual quality further improves, achieving near-GT or even better performance on the reconstruction setting. This indicates that jointly training the LBM and the decoder together breaks through the VAE reconstruction upper-bound.
Table 13: Evaluation results on a subset of VoiceFixer-GSR (Liu et al., 2022) with different sampling steps of VoiceBridge
Steps WV-MOS ($\uparrow$) UTMOS ($\uparrow$) DNSMOS ($\uparrow$) NISQA ($\uparrow$) PESQ ($\uparrow$)
steps=1 4.011 3.965 2.905 3.846 2.378
steps=2 4.105 4.062 3.005 3.924 2.522
steps=3 4.182 4.364 3.158 4.156 2.589
steps=4 4.291 4.367 3.205 4.160 2.692
steps=5 4.269 4.354 3.193 4.217 2.545
steps=10 4.201 4.358 3.200 4.296 2.445
steps=20 4.189 4.329 3.179 4.188 2.365
steps=50 4.145 4.330 3.161 4.269 2.270
Appendix H Inference Results with Different Sampling Steps
We test our GSR inference efficiency on a subset of VoiceFixer-GSR (Liu et al., 2022) with different sampling steps. As shown in Table 13, the best restoration results are achieved with around 4 sampling steps, similar to the observation in (Jukić et al., 2024). In the field of SE, a large number of sampling steps may not lead to better results, which has been observed in both diffusion and bridge-based models. In our analysis, one of the reasons is that, as the prior or condition information (namely the LQ observation) has already been very indicative of the HQ target, diffusion or bridge models are not required to generate the target with numerous iterative sampling steps. Given a large number of sampling steps, some of the sampling iterations may introduce error and result in sub-optimal generation quality. Meanwhile, we compare the inference efficiency of VoiceBridge with other GSR models, as shown in Table 14. VoiceBridge achieves the lowest Real-Time-Factor (RTF), defined as the time needed for restoring audio of unit time, among iterative models. Such a low RTF indicates that VoiceBridge is capable for real-time GSR systems. Experiments are conducted on a single A800 GPU.
@@ Appendix H Inference Results with Different Sampling Steps
RTF ($\downarrow$) 0.015 0.833 0.075 0.068
NFE 1 64 8 4
Table 15: Evaluation results on VoiceFixer-GSR (Liu et al., 2022) for different perceptual-aware fine-tuning processes of VoiceBridge.
Voicefixer-GSR (Liu et al., 2022) Finetuned Module Finetuning Objective PESQ ($\uparrow$) DNSMOS ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$) Bridge $\mathcal{L}_{\text{lat.}}$ 2.067 3.031 3.645 4.153 4.067 $\mathcal{L}_{\text{rec}}$ 2.125 3.049 3.649 4.207 3.524 $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gan}}$ 2.058 3.069 3.533 4.055 3.362 $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{hf}}$ 2.171 3.038 3.625 4.122 3.466 $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gan}}+\mathcal{L}_{\text{hf}}$ 2.182 3.049 3.599 4.158 3.487 Bridge ++ Decoder $\mathcal{L}_{\text{rec}}$ 2.297 2.952 3.588 3.972 2.301 $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gan}}$ 2.169 2.938 3.475 3.709 3.091 $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{hf}}$ 3.014 3.025 4.267 4.252 3.167 $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gan}}+\mathcal{L}_{\text{hf}}$ 2.722 3.091 4.228 4.366 4.172
@@ I.1 Experimental Setup
We conduct ablation studies on each innovation of VoiceBridge: EP-VAE, joint neural prior, and perceptual-aware post-training. For the first part of the ablation study, which mainly focuses on the variation of VAE networks, the VAE networks (with or without EP) are pre-trained for 500k steps on 8 A800 GPUs with a batch size of 16. The fine-tuning for joint neural prior costs 150k steps with the same settings. For fair comparison, the groups without joint neural prior are trained using the original objective for the same amount of steps. For the second part, which studies the different loss terms in the post-training stage, all variants are fine-tuned from the pre-trained bridge model in the main experiment for 50k steps on 8 A800 GPUs with a batch size of 16.
I.2 Further Details
Due to page limit, we show only partial results of the ablation study on the fine-tuning strategy and objective during the perceptual-aware finetuning stage in the main text. More results are demonstrated in Table 15. Based on the pre-trained LBM, we further conducted 9 groups of experiments with different settings in the post-training stage, to study the effect of different training objective and whether to adjust the decoder in this stage.
The experiments are divided into 2 main groups. In the first group, we keep the decoder frozen, and use the data-space loss between decoded audio and clean audio to fine-tune the bridge model. In the second group, the decoder and bridge are jointly fine-tuned together.
We test the following set of loss combinations, shown in Table 15. Firstly, a setting for continually training with the original objective (latent space MSE loss) $\mathcal{L}_{\text{latent}}$ is provided as baseline reference (Line 1). For the other settings, we include the MRSTFT reconstruction loss $\mathcal{L}_{\text{rec}}$, which is used for VAE training, as the basic data-space loss for the post-training stage (Line 2,6). Then, we add the GAN loss $\mathcal{L}_{\text{gan}}$, which is composed of the adversarial and feature matching loss and the human feedback term $\mathcal{L}_{\text{hf}}$ to the training objectives, first separately (Line 3,4,7,8), then together (Line 5,9). All these variants are applied to both the two groups in which the decoder is frozen or jointly finetuned. All hyper-parameters are aligned with the VAE training stage in Appendix D.
As shown in Table 15, conducting post-training with the decoder frozen makes little improvement to the restoration system, regardless of the training objective we use. Within the group which we do joint post-training of both the bridge and decoder, adding only the reconstruction loss or with the GAN loss for the decoder also has negligible or even negative effects. Interestingly, adding the human feedback term $\mathcal{L}_{\text{hf}}$ has a significant impact on the corresponding metrics, PESQ and UTMOS, but causes a decrease in other metrics. In fact, the generated audio in this group shares a common artifact, which severely degrades the listening experience. This phenomenon can be comprehended as that, because the scoring network or algorithm for metric evaluation is frozen during training, the decoder learns to attack the evaluation network or algorithm, generating special cases that hacks the perceptual evaluator. Adding the GAN loss and the human feedback term together successfully avoids such hacking, since the discriminator can easily spotify such hacking features. As shown in the last line of Table 15, adding all objectives achieves a consistent and signficant improvement in all metrics, demonstrating a real advancement in perceptual quality.

### v3

@@ G.1 Evaluation datasets
As GSR is defined as general restoration, including different restoration processes, the restoration process from a single degradation naturally becomes its subtask. To evaluate VoiceBridge’s zero-shot ability on these restoration subtasks, we choose the traditional Speech Enhancement (SE) task as a typical example. We utilize VoiceBank-Demand (Valentini-Botinhao, 2017) and WSJ0-CHiME3 (Barker et al., 2015). The VoiceBank-Demand dataset mixes utterances from the Voicebank speech corpus with real-world noise from the DEMAND dataset, serving as a benchmark testset for the SE. WSJ0-CHiME3, similarly, mixes clean speech signals from the WSJ0 dataset with noise from CHiME3, but with a smaller signal-to-noise ratio, making the denoising task harder. The two benchmark datasets are widely used in prior SE works (Lu et al., 2021, 2022; Richter et al., 2023; Lemercier et al., 2023b; Jukić et al., 2024).
@@
For the CAR test, we utilize the VCTK (Veaux and Yamagishi, 2017) testset, using Encodec (Défossez et al., 2022) to compress the 48kHz audio to a codec at a 3kbps frame rate, and then reconstruct it.
@@
Table 8: Denoising results on VCTK (Veaux and Yamagishi, 2017) with additive noise from DEMAND (Thiemann et al., 2013) at different SNR levels.
Dataset Method PESQ($\uparrow$) WVMOS($\uparrow$) UTMOS($\uparrow$) DNSMOS($\uparrow$) NISQA($\uparrow$)
SNR=-15dB SBSE 1.119 -0.949 1.480 1.513 1.869
Voicefixer 1.225 2.423 2.278 2.313 2.848
Resemble-Enhance 1.168 3.394 2.440 2.640 3.130
UniverSE++ 1.211 2.772 2.271 2.374 2.950
VoiceBrigde 1.352 3.479 2.755 2.698 3.270
SNR=-10dB SBSE 1.224 0.347 1.829 1.822 2.389
Voicefixer 1.588 3.263 2.845 2.791 3.576
Resemble-Enhance 1.345 3.783 2.929 2.876 3.693
UniverSE++ 1.527 3.704 3.075 2.933 3.971
VoiceBrigde 1.453 3.908 3.339 2.944 3.725
SNR=-5dB SBSE 1.410 1.698 2.301 2.275 2.933
Voicefixer 1.871 3.901 3.346 3.106 4.133
Resemble-Enhance 1.586 4.001 3.266 3.009 4.075
UniverSE++ 1.790 3.820 3.320 2.913 4.177
VoiceBrigde 1.730 4.093 3.641 3.022 3.969
SNR=0dB SBSE 1.663 2.938 2.826 2.665 3.468
Voicefixer 2.143 4.074 3.548 3.170 4.350
Resemble-Enhance 1.891 4.135 3.491 3.070 4.322
UniverSE++ 2.240 4.100 3.674 2.974 4.291
VoiceBrigde 2.200 4.200 3.765 3.098 4.201
G.2 Baseline Methods
G.3 Metrics
G.4 Degradation Breakdown
To make comprehensive evaluation, we include task-specific baseline models and evaluation metrics. For the BWE task, we include audio super-resolution models and other speech BWE model, including AudioSR (Liu et al., 2024), NuWave2 (Han and Lee, 2022), TFiLM (Birnbaum et al., 2019), and HiFi++ (Andreev et al., 2023). For dereverberation on WSJ-Reverb, we include SGMSE+ (Richter et al., 2023; Welker et al., 2022) and StoRM (Lemercier et al., 2023b). For SE, we include SBSE (Jukić et al., 2024). Additional metrics involves the following: STOI (Taal et al., 2010) is an objective metric that measures the intelligibility of degraded speech. LSD (Liu et al., 2022) directly calculates the log-spectrogram-distance between the enhanced and ground truth signals, which is a commonly used metric in BWE scenarios (Liu et al., 2024; Li et al., 2025a). SRMR (Falk et al., 2010), a non-intrusive metric specially designed for measuring signal to reverberation energy ratio, is utilized for measuring dereverberation performance in our experiments.
@@ G.4 Degradation Breakdown
VoiceBridge 3.47 4.03 3.19 4.50
G.5 Comparison with Pretrained Models
@@
G.6 Comparison with other Closed-Source Models
Table 10: Comparison with FINALLY (Babaev et al., 2024) and HiFi-GAN-2 (Su et al., 2021), with samples from FINALLY’s demo page.
FINALLY (Babaev et al., 2024) Demonstration Samples
Model DNSMOS ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$)
HiFi-GAN-2 3.283 3.947 3.470
FINALLY 3.313 3.847 3.989
VoiceBridge 3.093 4.122 4.061
Moreover, we compare VoiceBridge to other closed-source models such as FINALLY (Babaev et al., 2024), HiFi-GAN-2 (Su et al., 2021), SpeechFlow (Liu et al., 2023a), and Ku et al. (Ku et al., 2025). For comparison with FINALLY and HIFI-GAN-2, we use the open samples on the demonstration page of FINALLY at https://mmacosha.github.io/finally-demo/. The results are shown in Table 10. We can see that, VoiceBridge beat FINALLY and HiFi-GAN-2 on the WV-MOS and UTMOS metric, while reaching on-par performance on the DNSMOS metric, demonstrating competitive performance against strong closed-source baselines. We need to note that, FINALLY is pretrained on large internal datasets, where as VoiceBridge is only trained on publicly available data, which highlights the advantage of our modeling. Meanwhile, the testing samples are from the FINALLY’s demo page, which might be cherry-picked results for the baseline model, yet completely random for VoiceBridge.
Table 11: Comparison with SpeechFlow (Liu et al., 2023a) and Ku et al. (Ku et al., 2025), on corresponding benchmarks.
Downstream Task Evaluation on Curated Testsets
VB-Demand
Model PESQ ($\uparrow$) WVMOS ($\uparrow$) ESTOI ($\uparrow$) SISDR ($\uparrow$)
SpeechFlow 3.13 - 0.87 -
Ku et al. 3.27 4.41 0.88 19.1
VoiceBridge 2.83 4.39 0.79 4.8
WSJ0-CHiME3
Model PESQ ($\uparrow$) WVMOS ($\uparrow$) ESTOI ($\uparrow$) SISDR ($\uparrow$)
Ku et al. 2.85 4.26 0.92 16.2
VoiceBridge 1.74 4.37 0.83 3.4
WSJ0-BWE
Model WVMOS ($\uparrow$) SQUIM-MOS ($\uparrow$) ESTOI ($\uparrow$) SISDR ($\uparrow$)
Ku et al. 3.93 4.41 0.92 8.5
VoiceBridge 4.29 4.49 0.86 3.5
To compare against SpeechFlow and Ku et al.’s model, we evaluated VoiceBridge on the benchmark and metrics used in the corresponding papers. We reported the evaluation results on two denoising benchmarks VB-Demand (Valentini-Botinhao, 2017) and WSJ0-CHiME3 (Barker et al., 2015), and the bandwidth extension benchmark WSJ0-BWE (Ku et al., 2025). The results are shown in Table 11. VoiceBridge achievew comparable performance with the two baseline models on non-intrusive metrics, such as WVMOS and SQUIM-MOS, but showcase lower performance on other metrics. For this, we must state that: SpeechFlow and Ku et al.’s methods need specific finetuning for each task, while VoiceBridge is a general model without any task-specific adaptation. Meanwhile, both SpeechFlow and Ku et al.’s models work on 16kHz audio, while VoiceBridge generates 48kHz full-band audio, which brings significantly more challenges for the increased data dimension and high-frequency modeling. The used metrics only evaluate the lower 16kHz band, bringing inherent disadvantages to VoiceBridge. On waveform-level metrics like SISDR, latent-space methods like VoiceBridge typically underperforms data-level methods due to the latent reconstrucion error (Especially, VAE reconstruction is trained mainly on MR-STFT maginitude, not on the waveform dimension, and no phase preservation is guaranteed) and the cascading error from latent modeling. However, these metrics do not necessarily reflect the generated audio quality, as pointed out in (Manocha et al., 2022; Manjunath, 2009; Andreev et al., 2023; Babaev et al., 2024).
Appendix H Further Studies on Latent Modeling
H.1 Quantitative Results on the Joint Neural Prior
To quantitatively analyze the effect of prior convergence beyond tSNE visualization, we calculate the 2-Wasserstein distance (Panaretos and Zemel, 2019) matrix over the latent distribution for different priors. We model the latent distribution for each degradation type as a Gaussian distribution with the sample mean and variance. The 2-Wasserstein (Fréchet) distance between two Gaussian-fitted distributions $\mathcal{N}(\mu_{i},\Sigma_{i})$ and $\mathcal{N}(\mu_{j},\Sigma_{j})$ is defined as
@@ H.1 Quantitative Results on the Joint Neural Prior
H.2 Exceeding the Perceptual Quality Limits of Latent Models
Without our proposed post-training process with perceptual awareness, the LBM is trained to capture the distribution of the latent of the HQ target, which may not be fully aligned with human perceptual quality. Moreover, without joint post-training, the perceptual generation quality of VoiceBridge will be limited by VAE as well. As the perceptual objectives have not been explicitly considered in the training of VAE, the perceptual quality of VoiceBridge will be inevitably restricted. By the perceptual joint post-training process, we can improve the perceptual generation quality of both LBMs and the VAE decoder, meanwhile reducing their cascading error in generation.
@@ H.2 Exceeding the Perceptual Quality Limits of Latent Models
To verify that VoiceBridge surpasses the upper-bound reconstruction performance of the original VAE, we compare the restoration results of noisy inputs with EP-VAE reconstructions of clean ones. Specifically, we report four groups of results: (1) the EP-VAE reconstruction of clean audio (VAE-Rec.); (2) the restoration results of corresponding degraded audio of VoiceBridge (VB-Gen.); (3) the restoration results of VoiceBridge when passed the clean audio as input (VB-Rec.), which can be seen as a version of reconstruction with enhanced perceptual quality by VoiceBridge; (4) the ground-truth clean target (GT). Clean and degraded data are taken from the VoiceBank-Demand (Valentini-Botinhao, 2017) dataset. As shown in Table 12, the post-training stage enables VoiceBridge’s restoration to surpass the VAE reconstruction limits in perceptual quality. Moreover, when GT is passed as input for VoiceBridge instead of its noisy version, the perceptual quality further improves, achieving near-GT or even better performance on the reconstruction setting. This indicates that jointly training the LBM and the decoder together breaks through the VAE reconstruction upper-bound.
Table 13: VAE-only restoration results on different degradations
Curated Validation Sets for Different Degradations
Model PESQ ($\uparrow$) DNSMOS ($\uparrow$) WVMOS ($\uparrow$) UTMOS ($\uparrow$) NISQA ($\uparrow$)
Noise
VoiceBridge 2.34 3.17 4.30 4.26 4.18
VAE 1.71 2.58 3.51 3.37 3.23
Reverb
VoiceBridge 2.67 3.24 4.41 4.31 4.63
VAE 1.52 2.72 3.85 3.69 4.20
BWE
VoiceBridge 3.32 3.23 4.41 4.31 4.64
VAE 2.63 2.83 3.75 3.79 4.41
CLIP
VoiceBridge 3.57 3.23 4.43 4.31 4.69
VAE 2.94 2.82 3.96 3.90 4.44
H.3 VAE-only Restoration
In the joint neural prior training stage, the VAE is trained to align the degraded prior to clean prior in multiple measures, including those measuring reconstruction distances, which means the reconstruction of degraded speech will move forward the direction of the restoration target. After the post-training stage, the decoder is further adapted to generate speech with high perceptual quality. This indicate that the VAE itself have certain amount of restoration abilities. To further evaluate the ability of the VAE and the role of the latent bridge model, we test the performance of VAE-only restoration compared with VoiceBridge on different degradation types. The results are shown in Table 13. Basically, we can find that VAE-only enhancement works well on Bandwidth extension task and de-clipping tasks. These two tasks are considered easier, since they preserve the main structure of the speech signal. For the other harder tasks, generative modeling gives more effort. The bridge model successfully refines all degradations to a similar HQ leel beyond the VAE reconstruction.
@@ H.3 VAE-only Restoration
RTF ($\downarrow$) 0.015 0.833 0.075 0.025
NFE 1 64 8 1
@@ I.1 Experimental Setup
We conduct ablation studies on each innovation of VoiceBridge: EP-VAE, joint neural prior, and denoiser-to-generator post-training. For the first part of the ablation study, which mainly focuses on the variation of VAE networks, the VAE networks (with or without EP) are pre-trained for 500k steps on 8 A800 GPUs with a batch size of 16. The fine-tuning for joint neural prior costs 150k steps with the same settings. For fair comparison, the groups without joint neural prior are trained using the original objective for the same amount of steps. For the second part, which studies the different loss terms in the post-training stage, all variants are fine-tuned from the pre-trained bridge model in the main experiment for 50k steps on 8 A800 GPUs with a batch size of 16.
Table 15: Evaluation results on VoiceFixer-GSR (Liu et al., 2022) and DNS-Real (Reddy et al., 2021a) for bridge modeling in different spaces.
Voicefixer-GSR (Liu et al., 2022)
Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$)
Waveform SB 1.56 2.65 3.83 2.41 2.21 2.87 1.73
STFT SB 2.06 3.07 3.34 2.58 2.69 2.44 2.73
Latent SB 2.07 3.43 4.09 3.20 3.58 3.86 3.76
DNS-Real (Reddy et al., 2021a)
Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$)
Waveform SB - 2.64 2.55 1.94 1.61 1.55 1.56
STFT SB - 2.96 3.57 2.55 2.27 2.59 3.39
Latent SB - 3.39 4.03 3.12 3.15 3.69 4.34
I.2 Ablation on Bridge Modeling Space
To validate the necessity of latent modeling for the GSR task, we further conduct an ablation study on the modeling space for SB models. We train three bridge models on VAE latent, waveform, and complex STFT data space respectively. For the waveform-space bridge, we adopt the diffwave (Kong et al., 2020) architecture. For the STFT space, the NeMo (Harper et al., ) (adopted in  (Liu et al., 2023a) and  (Ku et al., 2025)) architecture is used. All networks are scaled to around 540M parameters by increasing the number of layers and hidden channels. For the VAE latent, we use VoiceBridge’s configuration without the denoiser-to-generator post-training, for equal comparison on the modeling spaces. All models are trained for 500k steps on the same data mixture as VoiceBridge. We evaluate the three models on the VoiceFixer-GSR and the DNS-Real benchmarks. The results are shown in Table 15 We can clearly observe that latent bridge models far outperform waveform and STFT domain bridge models, given the high information density of the compressed latent space. Meanwhile, the latent SB models take much less computation than the other two models, as the sequence modeled is much shorter.
Table 16: Evaluation results on VoiceFixer-GSR (Liu et al., 2022) and DNS-Real (Reddy et al., 2021a) for different joint neural prior training objectives
Voicefixer-GSR (Liu et al., 2022)
Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$)
KL-only 2.08 3.15 3.91 3.13 3.12 3.37 3.68
JNP 2.15 3.43 4.09 3.20 3.58 3.86 3.76
DNS-Real (Reddy et al., 2021a)
Model SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$)
KL-only 3.03 3.88 2.97 2.75 3.31 4.16
JNP 3.39 4.03 3.12 3.15 3.69 4.34
Table 17: Evaluation results on curated testsets with different degradation types for bridge models with and without joint neural prior.
Effect of JNP on Perceptual Metrics
Model PESQ ($\uparrow$) DNSMOS ($\uparrow$) WVMOS ($\uparrow$) UTMOS ($\uparrow$) NISQA ($\uparrow$)
Noise
w/ JNP 1.88 3.17 4.05 3.69 4.36
w/o JNP 1.75 2.95 4.01 3.65 4.10
Reverb
w/ JNP 2.12 3.24 4.11 3.77 4.60
w/o JNP 2.01 3.13 4.10 3.72 4.43
BWE
w/ JNP 2.70 3.30 4.32 3.97 4.69
w/o JNP 2.51 3.19 4.26 3.95 4.50
CLIP
w/ JNP 3.26 3.25 4.31 4.01 4.66
w/o JNP 2.95 3.16 4.20 3.91 4.63
I.3 Ablation on joint neural prior training objective design
A core difference of our joint neural prior with (Fang et al., 2021) is that we align the LQ and HQ priors not only by the Euclidean measure in latent space (wich is basically equivalent with the KL alignment done in (Fang et al., 2021) if regardless of the variance). In contrast, we add objectives including cosine similarity in latent space and multiple supervision signals in the reconstructed data space, in order to reduce the distance between the latents in all geometrical measures. We conduct an ablation study with two bridge models trained with the multi-measure aligned joint neural prior and KL-aligned-only prior respectively, to prove that our multi-measure alignment is necessary. The results are shown in Table 16. Our joint neural prior clearly outperforms the KL-only alignment proposed in (Fang et al., 2021), proving our method’s effectiveness.
I.4 Ablation on joint neural prior’s effectiveness for different degradation types
To further discuss whether the joint neural prior can converge the prior of diverse degradation types to a unified clean distribution, rather than only working on certain types, we conduct an ablation study on the effectiveness of joint neural prior for different degradation types. We train two bridge models with and without joint neural prior respectively, and evaluate their performance on different degradation types on curated testsets with additional noise, reverberation, bandwidth limitation, and clipping on the VCTK corpus. The models are trained for 500k steps. Results are shown in Table 17. The results demonstrate consistent improvement across all dereverberation types, indicating that JNP benefit all distortions simultaneously.